OpenStudy (anonymous):

Take the integral of sqrt(64-x^2) dx.

OpenStudy (anonymous):

u=64-x^2?

OpenStudy (ipwnbunnies):

Yes

OpenStudy (ipwnbunnies):

Oh, no, this is not a u-sub problem. Didn't look at your problem that well.

OpenStudy (anonymous):

I just knew it, then what's the strategy?

OpenStudy (ipwnbunnies):

Bleh, trig sub. Which I'm not fond of, or good at. u = sqrt(64)sin(x), du/dx = sqrt(64)cos(x). Not too good at the next part...

OpenStudy (ipwnbunnies):

Replace x with sqrt(64)sin(x) I mean. The terms underneath the square root total 64cos^2(x)

OpenStudy (ipwnbunnies):

Man, I'm losing it. It's something around those lines. Sorry for getting your hopes up lol.

OpenStudy (anonymous):

use trigonometric substitution first take factor from the sqrt

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{64-x^2}dx\] \[8\int\limits_{}^{}\sqrt{1-\frac{ x ^{2} }{ 64 }} dx\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

\[8\int\limits_{}^{}\sqrt{1-(\frac{ x }{ 8 })^{2}} dx\]

OpenStudy (anonymous):

np but can u complete it ?

OpenStudy (anonymous):

@Idealist